You have 30L of a 60% acid solution, how many liters of acid do you need to add to have a solution that has a concentration of 80%?
Answer by Earlsdon(6294)

Let x = the required number of liters of acid to be added to the 30 liters of 60% acid solution to obtain 30+x liters of 80% acid solution.
You can set up the equation to solve for x like this:
30(0.6)+x = (30+x)(0.8) Simplify and solve for x.
18+x = 24+0.8x Subtract 0.8x from both sides.
18+0.2x = 24 Now subtract 18 from both sides.
0.2x = 6 Finally, divide both sides by 0.2
x = 30 liters.